NPV.  Miglietti Restaurants is looking at a project with the following forecasted​ sales: ​ first-year sales quantity of 33,000​, with an annual growth rate of 4.00​% over the next ten years. The sales price per unit will start at ​$40.00 and will grow at 2.00% per year. The production costs are expected to be 55​% of the current​ year's sales price. The manufacturing equipment to aid this project will have a total cost​ (including installation) of​$2 comma 500 comma 0002,500,000. It will be depreciated using​MACRS, LOADING... ​, and has a​ seven-year MACRS life classification. Fixed costs will be ​$350,000 per year. Miglietti Restaurants has a tax rate of 38​%. What is the operating cash flow for this project over these ten​ years? Find the NPV of the project for Miglietti Restaurants if the manufacturing equipment can be sold for ​$140,000 at the end of the​ ten-year project and the cost of capital for this project is 99​%.

1. What is the operating cash flow for this project in year​ 1?

2. What is the operating cash flow for this project in year​ 2?

3. What is the operating cash flow for this project in year​ 3?

4. What is the operating cash flow for this project in year​ 4?

5. What is the operating cash flow for this project in year​ 5?

6. What is the operating cash flow for this project in year​ 6?

7. What is the operating cash flow for this project in year​ 7?

8. What is the operating cash flow for this project in year​ 8?

9. What is the operating cash flow for this project in year​ 9?

10. What is the operating cash flow for this project in year​ 10?

11. What is the after-tax cash flow of the project at disposal?

12. What is the NPV of the project?
Answer #1

Operating Cash Flow in year 1 is $287,035

Operating Cash Flow in year 2 is $406,326.42

Operating Cash Flow in year 3 is $363,579.25

Operating Cash Flow in year 4 is $341,276.24

Operating Cash Flow in year 5 is $331,185.21

Operating Cash Flow in year 6 is $362,444.31

Operating Cash Flow in year 7 is $392,617.33

Operating Cash Flow in year 8 is $382,059.09

Operating Cash Flow in year 9 is $373,535.79

Operating Cash Flow in year 10 is $409,440.37

Terminal Cash Flow in year 10 is $86,800

NPV of the Project is -$150,717.33 at 9%

NPV of the project is -$2,162,710.96 at 99%

Calculation of NPV of the Project
Particulars 0 1 2 3 4 5 6 7 8 9 10
Initial Investment
Manufacturing Equioment Cost (A) -2500000
Operating Cash Flows
Depreciation Rates (B) 14.29% 24.49% 17.49% 12.49% 8.93% 8.92% 8.93% 4.46%
Sales Quantity (B) 33000 34320 35692.8 37120.512 38605.33248 40149.54578 41755.52761 43425.74871 45162.77866 46969.28981
Annual Sales D = C* $40* (1.02)^n
n=0,1,2,3,4,5,6,7,8,9
1320000 1400256 1485391.565 1575703.372 1671506.137 1773133.71 1880940.24 1995301.406 2116615.732 2245305.968
Production Costs (E = D*55%) 726000 770140.8 816965.3606 866636.8546 919328.3753 975223.5405 1034517.132 1097415.773 1164138.652 1234918.283
Fixed Costs (F) 350000 350000 350000 350000 350000 350000 350000 350000 350000 350000
Depreciation (G = $2,500,000 * B) 357250 612250 437250 312250 223250 223000 223250 111500 0 0
Profit Before Tax (H = D-E-F-G) -113250 -332134.8 -118823.7958 46816.51737 178927.7616 224910.1695 273173.1078 436385.6328 602477.0793 660387.6857
Tax @38% (I = H*38%) -43035 -126211.224 -45153.04242 17790.2766 67992.54942 85465.86442 103805.781 165826.5405 228941.2901 250947.3206
Profit After Tax (J = H-I) -70215 -205923.576 -73670.75342 29026.24077 110935.2122 139444.3051 169367.3269 270559.0923 373535.7892 409440.3651
Add back Depreciation (K = G) 357250 612250 437250 312250 223250 223000 223250 111500 0 0
Net Operating Cash Flows (L = J+K) 287035 406326.424 363579.2466 341276.2408 334185.2122 362444.3051 392617.3269 382059.0923 373535.7892 409440.3651
Terminal Value
Sale Value (M) 140000
Tax @38% (N = M*38%) 53200
After tax sale value (O = M-N) 86800
Total Cash Flows (P = A+L+O) -2500000 287035 406326.424 363579.2466 341276.2408 334185.2122 362444.3051 392617.3269 382059.0923 373535.7892